• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name kelleher

    Log entry time 16:07:28 on July11,2010

    Entry number 331479

    keyword=Day Shift Summary -- June 11, 2010

    Sunday, July 11 Day Shift

    Plan: production data

    08:00 Start shift in production mode, 120uA on Ta target, run #2044
    09:40 Receiving 100uA beam. MCC says that they are getting trips from their beam loss monitor. They were worried that somethingwas activated and tried to run at lower current (100uA).
    10:30 returned to quasi-stable 120uA current
    10:57 Called MCC to confirm that beam is visible in dump viewer
    13:15 Shift checklist
    16:00 End shift during run #2055


    Run Summary

    Run Begin End Notes
    -----------------------------------
    2044 07:33 08:17 499k
    2045 08:18 09:03 499k
    2046 09:04 09:56 499k
    2047 09:56 10:15 195k no beam for last 15 minutes (running at 100uA max)
    2048 10:32 11:24 499k
    2049 11:26 12:11 499k
    2050 12:12 12:57 499k
    2051 12:58 13:39 499k
    2052 13:40 14:05 265k ROC1 was not responding
    2053 14:08 14:49 499k No beam trips after first 5 minutes
    2054 14:50 15:32 499k 2 beam trips over the entire run
    2055 15:32 ended shift taking data